OpenStudy (anonymous):

Give the sum of this arithmetic series: 2+4+6+...+100

OpenStudy (lgbasallote):

first, find out how many terms there are by doing \[a_n = a_ 1 + (n-1)d\] an is the last term in this case 100 a1 is the first term, in this case 2 d is the common difference and n is what you're looking for

OpenStudy (lgbasallote):

can you try doing it?

OpenStudy (lgbasallote):

solve for n

OpenStudy (cwrw238):

sum = (n/2) [ a + l]

Parth (parthkohli):

\[{n \over 2}\left(a_1 + a_n \right)\]First, find the number of terms.\[a_n = a_1 + (n - 1)d\]\[a_n = 2 + 2(n - 1)\]solve for n

OpenStudy (anonymous):

161 terms, right?

Parth (parthkohli):

lol no

OpenStudy (anonymous):

Oh I see what i did, give me one sec man

OpenStudy (lgbasallote):

sure

OpenStudy (lgbasallote):

take yur time

OpenStudy (cwrw238):

number of terms n = 50 a = 2 and l last term = 100

OpenStudy (anonymous):

Can you show your work on how you got the 50 for n?

OpenStudy (lgbasallote):

if you substitute the values \[\implies 100 = 2 + (n-1)2\] you get that right?

OpenStudy (anonymous):

Right the common difference is 2

OpenStudy (lgbasallote):

right. so if you distribute 2 into n-1 you get \[\implies 100 = 2 + 2n - 2\] right?

OpenStudy (anonymous):

so 100=2n n=50

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

Then we need to use the sum of an aritmetic series formula, correct?

OpenStudy (lgbasallote):

so now, you can use the formula \[S_n = \frac{n(a_1 + a_n)}{2}\]

OpenStudy (lgbasallote):

yes you're right

OpenStudy (anonymous):

Thank you. Hold on real quick as I complete the problem

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

So you would set it up as this, correct: s(sub 50)=50(2+100)/2

OpenStudy (lgbasallote):

yes!

OpenStudy (anonymous):

And get (25)(102), and get 2550?

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

Can also think of it like this: 2+4+6+8+...+100 = (2+100)+(4+98)+(6+96)+(8+94)+...(50+52). If you can't remember the formula, this is how you derive it.

OpenStudy (anonymous):

Interesting, thanks guys Hold one, I have one more to try

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

Ok, the next problem is 1+5+9+...+401 So, first thing I did was use the arithmetic sequence formula a to the nth term=a1+(n-1)d So i plugged in 401=1+(4n-4) Ended up getting 101 for n Then I used the sum of an aritmetic series formula s(101)=101(1+401)/2 Ended up with 40602/2 And got 20301 as the sum of all 101 terms. Is that correct Igbasallote?